# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ V-D4 Symmetry-based near-field estimation algorithms
Another strategy to reduce the computational complexity of the 2D peak search is by exploiting the symmetry of array structure. Here, we present two approaches to decouple the problems of angle and distance estimation by leveraging the symmetric structure of the antenna array for mutiple targets.
Modified MUSIC [134]: After computing the covariance matrix of the received signal ${\bf{R}}_{X}$ in (75), its anti-diagonal elements without noise are given by
@@
where $p_{k}$ denotes the equivalent transmitted signal power of target $k$, i.e., $\mathbb{E}\{{\mathbf{s}^{H}_{k}}{\mathbf{s}_{k}}\}=p_{k}$. From the anti-diagonal samples of the covariance samples we can reconstruct a new vector that depends only on the angle parameters, given by
$[\tilde{\mathbf{y}}]_{m}=\sum_{k=1}^{K}p_{k}e^{j2m{\omega_{k}}}.$ (91)
To perform 1D MUSIC, a full-rank set of observations is required, so spatial smoothing techniques are employed by splitting $\tilde{\mathbf{y}}$ into $L$ subvectors, each comprising $(M+1-L)$ entries. The $\ell$-th split signal can be expressed as
@@
where $\mathbf{R}_{p}=\sum_{\ell=1}^{L}\mathbf{p}_{\ell}\mathbf{p}_{\ell}^{H}$. To ensure that $\mathbf{R}_{p}$ is full-rank, we impose the constraint $K<L$. After obtaining the angles of targets, we can perform a 1D search in the distance domain.
TABLE VI: Performance comparison for near-field sensing algorithms
@@
Generalized ESPRIT [138]: For generalized ESPRIT, the antenna array is partitioned into two subarrays, with the first subarray comprising the first $J$ antennas arranged in an ascending order, and the second subarray consisting of the last $J$ antennas arranged in a descending order. The steering vectors of the two subarrays can be expressed as
@@
$\mathbf{a}_{2}$ $=\!\!\left[e^{j\left(N\omega_{k}+N^{2}\psi_{k}\right)},\cdots,c^{j\left((N-J+1)\omega_{k}+(N-J+1)^{2}\psi_{k}\right)}\right]^{\mathrm{T}}.$
@@
$=\mathbf{D}(\theta)\mathbf{A}_{1},$
when $\theta$ aligns with the $i$-th target, and the $i$-th column of $\mathbf{A}_{3}$ is identical to the $i$-th column of $\mathbf{A}_{1}$. Matrix $\mathbf{F}(\theta)=\mathbf{J}\mathbf{E}_{s2}-\mathbf{D}(\theta)\mathbf{E}_{s1}$ is rank deficient when $\theta$ corresponds to one of the target angles, which means that for an arbitrary full column rank matrix $\mathbf{W}\in\mathbb{C}^{J\times K}$, matrix $\mathbf{W}^{H}\mathbf{F}(\theta_{k})$ is singular and its determinant is zero. The generalized ESPRIT exploits this observation by defining the following spectrum function whose peaks will correspond to the angles of targets
@@
Once the targets’ angles are estimated, a 1D search can be conducted to estimate the distance parameters.
@@ V-E Performance Analysis and Comparison
Table VI shows the properties of the near-field sensing algorithms discussed above, including their complexity and resolution capability. Fig. 16 gives a numerical representation of the theoretical complexity of each algorithm based on the values in Table VI for receive array sizes of $M_{r}=64,128$ and $512$. The search grids for the angle and distance parameters are chosen to have $n_{g}=\frac{180}{0.05}=3600$ and $n_{l}=\frac{90}{0.1}=900$ cells, respectively. It is observed that for all $M_{r}$, SoC and modified MUSIC have the lowest complexity, while 2D MUSIC and 2D BF have the highest. However, as $M_{r}$ increases, the complexity gap between 2D MUSIC/2D BF and other algorithms becomes smaller, because the complexity of the EVD becomes more dominant for large $M_{r}$.
Figs. 17 and 18 show the RMSE of the angle and distance parameters versus the SNR, respectively. The number of antenna elements is $M_{r}=64$ with $d=\frac{\lambda}{4}$. The target is located at $(10.2^{\circ},5.64\ \rm{m})$. As the SNR increases, the estimation accuracy of all algorithms improves. For angle estimation, RD and RR perform best, thanks to their high-sensitivity angle search procedures. On the other hand, SoC based algorithm performs worst since the exploited structure of the covariance only holds asymptotically. We further see that 2D BF and 2D MUSIC are sensitive to SNR, and can only achieve good estimation performance when the SNR exceeds $-5$ dB.
@@
In practice, several challenges still exist when implementing the existing algorithms as elaborated below.
$\bullet$ Element spacing requirement: To avoid phase ambiguity after calculating the second-order correlation for the SoC-based algorithm, the inter-antenna spacing should satisfy $d\leq\frac{\lambda}{4}$, which may lead to high mutual coupling between antenna and is not compatible with existing communication systems that typically have half-wavelength spacing.
$\bullet$ Reduced resolution: For the SoC algorithm, the dimension of the data vector in (79) is only $dM_{r}/2$ due to the process of constructing cumulants. Therefore, the size of the effective aperture is reduced, and the spatial resolution decreases from $\frac{2\lambda}{D}$ to $\frac{4\lambda}{D}$.
$\bullet$ Prior knowledge of the number of targets: Almost all the algorithms mentioned above require prior knowledge of the number of targets, so the ability to accurately obtain this information will greatly impact the estimation performance. For MUSIC, this can be achieved by calculating the number of significantly large eigenvalues. For RR and RD algorithms, the number of targets is crucial to properly choose the number of peaks in the angle spectrum.
$\bullet$ Error propagation: By decoupling the angle and distance parameters, the complexity by parameter search algorithms can be significantly reduced. However, the accuracy of distance parameter estimate is highly dependent on the estimation results for the angle parameter. For the extreme case when several targets are located in the same direction, SoC and RD algorithms are no longer applicable, which may lead to severe errors in the subsequent distance estimation.
$\bullet$ Parameter pairing: Parameter pairing is an additional step after the estimation of $K$ separate angles and distance values. For algorithms involving a 2D search, parameters are naturally paired. In addition, for the SoC-AoA and RR algorithms, the distances are obtained by performing a 1D search for each estimated angle, thus avoiding the problem of parameter pairing as well. For other algorithms, the problem and complexity associated with parameter pairing should be considered.
@@ VI Extended Discussion For MIMO-OFDM ISAC
In this section, we point out open problems and outline promising avenues that warrant further investigation for MIMO-OFDM ISAC to inspire future work.
@@ VI-A Wideband XL-MIMO OFDM ISAC
To achieve higher angular and delay resolution, larger scale MIMO and wider bandwidths are needed for MIMO-OFDM ISAC systems. However, wideband XL-MIMO OFDM ISAC introduces spatial wideband effects [139, 140], also known as the beam squint effects [141], where the propagation delay across the array aperture can not be ignored. The spatial wideband effect means that the array steering vector is dependent not only on the location or angle of the sensing target but also on frequency. As a result, the phase shift components in the spatial and subcarrier domains are no longer independent but become coupled for XL-MIMO OFDM ISAC. In addition, the use of XL-MIMO causes a variable Doppler frequency across the array aperture for mobile targets in the near-field, resulting in the coupling of the phase shift components in the spatial and Doppler domains. These challenging issues will fundamentally affect target sensing in XL-MIMO OFDM ISAC systems. Thus, the sensing methods and results established in this paper for conventional MIMO OFDM ISAC systems will require reformulation and modification.
@@ VI-B Signal Pre-processing for MIMO-OFDM ISAC
Clutter rejection and clustering: Clutter from environmental scatterers makes it difficult to distinguish sensing targets from other reflections. One efficient approach for clutter rejection is through the use of adaptive filtering methods, such as MMSE and Kalman filtering [142]. Additionally, techniques such as spatial filtering and Doppler processing can enhance the ability of the radar to distinguish between targets and clutter, especially in dynamic environments [143]. Clustering methods also play a significant role in improving the SNR and facilitating more accurate detection in complex sensing environments. For instance, K-means clustering and Gaussian mixture models (GMM) can be employed to effectively separate target echoes from clutter [144]. Emerging techniques based on machine learning are increasingly explored for clutter rejection in ISAC systems. For instance, convolutional neural networks (CNNs) have shown promise in automatically learning the features that distinguish clutter from desired signals, leading to improved classification and interference rejection [145]. Filtering clutter based on pre-stored channel information in CKM is also an emerging and promising approach for clutter rejection. As an example, the authors of [74] proposed a new CKM approach referred to as CLAM, which can effectively remove environmental clutter through pre-acquired clutter angle information.
@@
Synchronization: For bistatic and multi-static ISAC systems, the ISAC-TX and RX are spatially separated and not synchronized since they have separate LO and sampling clock references. Therefore, the received signal experiences an additional STO $\tau_{\Delta}$, CFO $f_{\Delta}$, and a sampling frequency offset (SFO) $\delta_{\Delta}$ caused by the deviation in the sampling frequency between the digital-to-analog converters (DAC) at the ISAC-TX and the analog-to-digital converters (ADC) at the RX[37]. Typically, synchronization is performed based on the LoS channel. Therefore, the RX needs to estimate and compensate for the time offset (TO) $\tau_{off}=\tau_{0}+\tau_{\Delta}$, the frequency offset (FO) $f_{off}=v_{0}+f_{\Delta}$, and $\delta_{\Delta}$. The TO must be compensated for first to determine the starting point for the OFDM signal. A coarse TO and FO can be estimated based on the C&S algorithm [146]. Furthermore, cross-correlation between the received signal and the first preamble symbol is performed for fine TO estimation[147]. After finding the starting point of the received signal, the SFO can be estimated based on the remaining preambles via a weighted LS algorithm [148]. For more accurate SFO estimation, the entire received signal can be interpolated by a multi-rate finite impulse response (FIR) filter [149], [150] and fed into a sample rate converter based on a polynomial filter [151]. Finally, based on the estimate of the coarse FO, the residual FO can be obtained by constructing a delay Doppler profile (DDP) based on pilots[152].
@@ VI-C Resource Allocation and Beam Management
In practice, an ISAC system cannot continuously occupy all time-frequency resources to execute a single communication or sensing task. For example, in 5G NR TDD mode, the slots for uplink and downlink communications are alternating and non-uniform [153, 154], presenting a new challenge for sensing tasks in ISAC systems. For practical multi-user ISAC scenarios, users obtain C&S services through orthogonal frequency division multiple access (OFDMA). Thus, different users can only be allocated part of the time-frequency resources and transmit beamforming is exploited to serve multiple users to accomplish different C&S tasks in ISAC systems. Therefore, resource allocation and beam management are two key design issues in ISAC systems, as discussed in the following.
@@
Resource allocation: In conventional communication systems, the collocated allocation of time-frequency resources inherently constrains the achievable delay and Doppler resolution for individual users. In ISAC systems, sparse resource allocation [153, 155] emerges as a promising alternative, enhancing sensing resolution by expanding the effective time and frequency spans. Sparse allocation can be broadly categorized into uniform and non-uniform approaches. Uniform sparse allocation employs fixed-interval resource distribution, which improves sensing resolution at the expense of a significantly reduced unambiguous range. Non-uniform sparse allocation maintains the unambiguous range while enhancing resolution, albeit at the cost of elevated sidelobes in delay/Doppler periodograms, potentially obscuring weak targets. Pre-storing non-uniform sparse patterns with suppressed sidelobes via traversal search and structured allocation based on nested[156], coprime[154], or minimum redundancy[157], coupled with differential coarray processing, are feasible non-uniform sparse allocation strategies. However, traversal search methods suffer from prohibitive computational complexity, while fixed allocation structures may conflict with functional reference signals in mobile communication standards. Consequently, critical challenges remain in optimizing time-frequency resources and subcarrier power allocation [40], necessitating further investigation into trade-offs between fairness and efficiency, as well as between C&S performance.
@@
Beam management: Beam management [158, 159] can be categorized as referring to beam alignment, beam sweeping, or beam tracking, based on different transmission or reception methods. Beam alignment involves aligning the transmission and reception beams between a transmitter and a receiver to maximize the signal transmission efficiency. Beam sweeping is the most commonly used method for achieving beam alignment, which involves transmitting or receiving beams from a predefined codebook over a given time interval to cover a given spatial area. Beam tracking aims to achieve beam alignment by predicting and tracking the positions or angles of moving users or targets. These various beam management techniques alter the transmit beamforming vectors for different time-frequency resources across different CPIs and affect the duration of a CPI. The signal processing methods discussed in this paper are applicable within a given CPI for each beam management technique. However, multiple beams under different management schemes may be transmitted or received simultaneously. Orthogonal multiple access (OMA) [160] can separate them via orthogonal time-frequency resources, whereas nonorthogonal multiple access (NOMA) [161] presents additional challenges. Furthermore, multi-beam transmission requires power allocation among the beams, which is another critical consideration.

### v2

@@ V-D4 Symmetry-based near-field estimation algorithms
Another strategy to reduce the computational complexity of the 2D peak search is by exploiting the symmetry of the array. Here, we present two approaches to decouple the problems of angle and distance estimation by leveraging the symmetric structure of the antenna array for multiple targets.
Modified MUSIC [150]: In the absence of noise, the anti-diagonal elements of the covariance ${\bf{R}}_{X}$ in (75) are given by
@@
where $p_{k}$ denotes the equivalent reflected signal power of target $k$, i.e., $\mathbb{E}\{{\mathbf{s}^{H}_{k}}{\mathbf{s}_{k}}\}=p_{k}$. From the anti-diagonal samples of the covariance, we can reconstruct a new vector that depends only on the angle parameters, given by
$\tilde{\mathbf{y}}]_{m}=\sum_{k=1}^{K}p_{k}e^{j2m{\omega_{k}}}.$ (91)
To perform 1D MUSIC, a full-rank set of observations is required, so spatial smoothing techniques are employed to split $\tilde{\mathbf{y}}$ into $L$ subvectors, each comprising (M+1L)(M+1-L) entries. The $\ell$-th split signal can be expressed as
@@
where $\mathbf{R}_{p}=\sum_{\ell=1}^{L}\mathbf{p}_{\ell}\mathbf{p}_{\ell}^{H}$. To ensure that $\mathbf{R}_{p}$ is full-rank, we impose the constraint $K<L$. After obtaining the target angles, we can perform a 1D search in the distance domain.
TABLE IX: Performance comparison for near-field sensing algorithms
@@
Generalized ESPRIT [154]: For generalized ESPRIT, the antenna array is partitioned into two subarrays, with the first subarray comprising the first $J$ antennas arranged in ascending order, and the second subarray consisting of the last $J$ antennas arranged in descending order. The steering vectors of the two subarrays can be expressed as
@@
$\mathbf{a}_{2}$ $=\!\!\left[e^{j\left(N\omega_{k}+N^{2}\psi_{k}\right)},\cdots,e^{j\left((N-J+1)\omega_{k}+(N-J+1)^{2}\psi_{k}\right)}\right]^{\mathrm{T}}.$
@@
$=\mathbf{D}(\theta)\mathbf{A}_{1}.$
When $\theta$ aligns with the $i$-th target, the $i$-th column of $\mathbf{A}_{3}$ is identical to the $i$-th column of $\mathbf{A}_{1}$, and $\mathbf{F}(\theta)=\mathbf{J}\mathbf{E}_{s2}-\mathbf{D}(\theta)\mathbf{E}_{s1}$ becomes rank deficient, which means that for any arbitrary full column rank matrix $\mathbf{W}\in\mathbb{C}^{J\times K}$, $\mathbf{W}^{H}\mathbf{F}(\theta_{k})$ is singular and its determinant is zero. For instance, $\mathbf{W}$ can be a semi-unitary matrix satisfying $\mathbf{W}^{H}\mathbf{W}=\mathbf{I}_{K}$. A simple example is
$\mathbf{W}=\begin{bmatrix}\mathbf{I}_{K}\\ \mathbf{0}_{(J-K)\times K}\end{bmatrix}$. (101)
The generalized ESPRIT algorithm exploits this observation by defining the following spectrum function whose peaks will correspond to the target angles
@@
Once the target angles are estimated, a 1D search can be conducted to estimate the distance parameters.
@@ V-E Performance Analysis and Comparison
Table IX shows the properties of the near-field sensing algorithms discussed above, including their complexity and resolution capability. Fig. 22 gives a numerical representation of the theoretical complexity of each algorithm based on the values in Table IX for receive array sizes $M_{r}=64,128$ and $512$. The search grids for the angle and distance parameters are chosen to have $n_{g}=\frac{180}{0.05}=3600$ and $n_{l}=\frac{90}{0.1}=900$ cells, respectively. It is observed that for all $M_{r}$, SoC and modified MUSIC have the lowest complexity, while 2D MUSIC and 2D BF have the highest. However, as $M_{r}$ increases, the complexity gap between 2D MUSIC/2D BF and other algorithms becomes smaller, because the complexity of the EVD becomes more dominant for large $M_{r}$.
Figs. 23 and 24 show the RMSE of the angle and distance parameters versus SNR, respectively. The number of antenna elements is $M_{r}=64$ with $d=\frac{\lambda}{4}$. The target is located at $(10.2^{\circ},5.64\ \rm{m})$. As the SNR increases, the estimation accuracy of all algorithms improves. For angle estimation, RD and RR perform best, thanks to their high-sensitivity angle search procedures. On the other hand, the SoC-based algorithm performs the worst since the exploited structure of the covariance holds asymptotically. We further see that 2D BF and 2D MUSIC are sensitive to SNR, and can only achieve good estimation performance when the SNR exceeds 5-5 dB.
@@
V-F Lessons Learned
Compared with the far-field case, the spherical wavefronts in near-field scenarios bring new opportunities and challenges for sensing. In particular, the nonlinear phase of the steering vector enables distance estimation without the need for propagation delay estimation, but the highly coupled parameters in the angle and distance domains also increase the estimation complexity. 2D algorithms and reduced dimension/rank algorithms achieve high RMSEs for angle and distance estimation, but also suffer from higher complexity. Modified MUSIC and SoC-based algorithms have lower complexity, but their RMSE performance is degraded. FFT-enhanced algorithms are able to achieve a good balance between performance and complexity. Several challenges still need to be addressed when implementing existing algorithms, as elaborated below.
  • Element spacing constraint: Most near-field sensing algorithms require the inter-element spacing to be smaller than a quarter wavelength. However, such dense array configurations may significantly increase the mutual coupling among antenna elements.
  • Degraded spatial resolution: The reduced element spacing inevitably shrinks the overall physical aperture, thereby reduing the inherent spatial resolution from $\frac{2\lambda}{D}$ to $\frac{4\lambda}{D}$. Moreover, for statistics-based algorithms such as the SoC method, the symmetric correlation operation combines each pair of antennas into a single output, which further halves the effective aperture and consequently results in an additional loss of spatial resolution.
  • Prior knowledge of the number of targets: Almost all the algorithms mentioned above require prior knowledge of the number of targets, so the ability to accurately obtain this information will greatly impact the estimation performance. For MUSIC, this can be achieved by calculating the number of significantly large eigenvalues. For RR and RD algorithms, the number of targets is crucial to properly choose the number of peaks in the angle spectrum.
  • Error propagation: By decoupling the angle and distance estimation, computational complexity can be significantly reduced. However, the distance parameter estimation accuracy is highly dependent on the quality of the angle estimates. For the extreme case when several targets are located in the same direction, SoC and RD algorithms are no longer applicable, precluding accurate distance estimation.
  • Parameter grouping: Parameter grouping is an additional step after separate estimation of angle and distance. However, algorithms involving a 2D search naturally lead to matched parameter estimates. For the SoC-AoA and RR algorithms, the distances are obtained by performing a 1D search for each estimated angle, thus the problem of parameter grouping is also avoided. For other algorithms, the problem and complexity associated with parameter grouping should be considered.
@@ VI Extended Discussion for MIMO-OFDM ISAC
In this section, we discuss open problems and outline promising avenues that warrant further investigation for MIMO-OFDM ISAC to inspire future work.
@@ VI-A Wideband XL-MIMO OFDM ISAC
Deploying a huge number of antennas and transmission in the mmWave/THz frequency bands can significantly enlarge the array aperture and exploit enormous bandwidth resources for MIMO-OFDM ISAC systems, thereby enhancing the resolution in both the spatial/angular and delay domains. However, wideband XL-MIMO OFDM ISAC introduces spatial wideband effects [155, 156], also known as beam squint and Doppler squint [157], where the propagation delay and Doppler spread across the array aperture cannot be ignored. Spatial wideband effects occur when the frequency dependence of the array response cannot be ignored. In such cases, the phase of the array response is a function of the subcarrier index. In addition, the use of extremely large arrays causes the target Doppler frequency to vary across the array aperture, resulting in the coupling of the phase shift components in the spatial and Doppler domains. These challenging issues will fundamentally affect target sensing in XL-MIMO OFDM ISAC systems. Thus, the sensing methods and results established in this paper for conventional systems will require reformulation and modification.
@@ VI-B Signal Pre-processing for MIMO-OFDM ISAC
Clutter rejection and clustering: Clutter from environmental scatterers makes it difficult to distinguish sensing targets from other reflections. An effective approach for clutter rejection is to use adaptive filtering methods, such as MMSE and Kalman filtering [158]. Additionally, techniques such as spatial filtering and Doppler processing can enhance the ability of the radar to distinguish between targets and clutter, especially in dynamic environments [159]. Clustering methods also play a significant role in improving the SNR and facilitating more accurate detection in complex sensing environments. For instance, K-means clustering and Gaussian mixture models (GMM) can be employed to effectively separate target echoes from clutter [160]. Emerging techniques based on machine learning are increasingly explored for clutter rejection in ISAC systems. For instance, convolutional neural networks (CNNs) have shown promise in automatically learning the features that distinguish clutter from desired signals, leading to improved classification and interference rejection [161]. Filtering clutter based on pre-stored channel information in CKM is also an emerging and promising approach for clutter rejection. As an example, the authors of [74] proposed a new CKM approach referred to as clutter angle map (CLAM) that can effectively remove environmental clutter through pre-acquired clutter angle information.
@@
Synchronization: For bistatic and multi-static ISAC systems, the ISAC-TX and RX are spatially separated and not synchronized since they have separate LO and sampling clock references. Therefore, the received signal experiences an additional STO $\tau_{\Delta}$, CFO $f_{\Delta}$, and a sampling frequency offset (SFO) $\delta_{\Delta}$ caused by the deviation in the sampling frequency between the DAC at the ISAC-TX and the ADC at the RX[37]. Typically, synchronization is performed based on the LoS channel. Therefore, the RX needs to estimate and compensate for the time offset (TO) $\tau_{off}=\tau_{0}+\tau_{\Delta}$, the frequency offset (FO) $f_{off}=v_{0}+f_{\Delta}$, and $\delta_{\Delta}$. The TO must be compensated for first to determine the starting point for the OFDM signal. A coarse TO and FO can be estimated based on C&S algorithms [162]. Furthermore, cross-correlation between the received signal and the first preamble symbol is performed for fine TO estimation [163]. After finding the starting point of the received signal, the SFO can be estimated based on the remaining preambles via a weighted LS algorithm [164]. For more accurate SFO estimation, the entire received signal can be interpolated by a multi-rate finite impulse response (FIR) filter [165], [166] and fed into a sample rate converter based on a polynomial filter [167]. Finally, the coarse FO estimate can be used to find the residual FO by constructing a delay Doppler profile (DDP) based on pilots[168].
@@ VI-C Resource Allocation and Beam Management
In practice, an ISAC system cannot continuously occupy all time-frequency resources to execute a single communication or sensing task. For example, in 5G NR TDD mode, the slots for uplink and downlink communications are alternating and non-uniform [169, 170], presenting a new challenge for sensing tasks in ISAC systems. For practical multi-user ISAC scenarios, users obtain C&S services through orthogonal frequency division multiple access (OFDMA). Thus, different users are allocated only part of the time-frequency resources and transmit beamforming is exploited to serve multiple users to accomplish different C&S tasks. Therefore, resource allocation and beam management are two key design issues in ISAC systems, as discussed in the following.
@@
Sparse resource allocation: In conventional communication systems, the collocated allocation of time-frequency resources inherently constrains the achievable delay and Doppler resolution for individual users. In ISAC systems, sparse resource allocation [169, 171] emerges as a promising alternative, enhancing sensing resolution by expanding the effective time and frequency spans. Sparse allocation can be broadly categorized into uniform and non-uniform approaches. Uniform sparse allocation employs fixed-interval resource distribution, which improves sensing resolution at the expense of a significantly reduced unambiguous range. Non-uniform sparse allocation maintains the unambiguous range while enhancing resolution, albeit at the cost of elevated sidelobes in delay/Doppler periodograms, potentially obscuring weak targets. Pre-storing non-uniform sparse patterns with suppressed sidelobes via traversal search, and structured allocation based on coprime [170], nested [172], or minimum redundancy [173], coupled with differential coarray processing, are feasible allocation strategies. However, traversal search methods suffer from prohibitive computational complexity, while fixed allocation structures may conflict with functional reference signals in mobile communication standards. Consequently, critical challenges remain in optimizing time-frequency resources and subcarrier power allocation [40], necessitating further investigation into trade-offs between fairness and efficiency, as well as between C&S performance.
Non-orthogonal multiple access: Non-orthogonal multiple access (NOMA) is another promising technology for enhancing the allocation of wireless resources in ISAC systems [174, 175]. By enabling the simultaneous provision of C&S services for multiple users within the same spatial, time, frequency, and code domains, NOMA can circumvent inherent limitations on C&S performance imposed by resource allocation in orthogonal multiple access (OMA). Current NOMA schemes primarily include power-domain [174] and code-domain [176] approaches. Power-domain NOMA allocates power through superposition coding (SC) and employs successive interference cancellation (SIC) to distinguish signals from different users, while code-domain NOMA relies on dedicated non-orthogonal sparse codebooks and message passing algorithms for user signal separation. However, the C&S performance of NOMA is constrained by the error propagation effect in SIC. Further research is warranted to model and mitigate residual inter-user or inter-C&S interference [177, 178].
@@
Beam management: Beam management [179, 180] can refer to beam alignment, beam sweeping, or beam tracking, based on different transmission or reception methods. Beam alignment involves aligning the beams between a transmitter and a receiver to maximize the signal transmission efficiency. Beam sweeping is the most commonly used method for achieving beam alignment, which involves transmitting or receiving beams from a predefined codebook over a given time interval to cover a given spatial area. Beam tracking aims to achieve beam alignment by predicting and tracking the positions or angles of moving users or targets. Beam management techniques alter the transmit beamforming vectors for different time-frequency resources across different CPIs and affect the duration of a CPI. The signal processing methods discussed in this paper are applicable within a given CPI for each beam management technique. However, multiple beams under different management schemes may be transmitted or received simultaneously. OMA can separate them via orthogonal time-frequency resources, whereas the use of NOMA presents additional challenges. Furthermore, multi-beam transmission requires power allocation among the beams, which is another critical consideration.